REASSIGNMENT OF MILLIMETERWAVE SPECTRUM OF THE HCN INTERNAL ROTATION BANDS OF H_{2}-HCN

KENSUKE HARADA, RISA YAMANAKA, and \underline{KEIICHI TANAKA}, Department of Chemistry, Faculty of Sciences, Kyushu University, Fukuoka, 812-8581 JAPAN.

The H_{2}-HCN complex is a weakly bound molecular complex and we have reported the pure rotational transitions of H_{2}-HCN in the MMW region.^{a} According to the results, ortho- and para-H_{2} complexes have different structures in the ground state, H_{2} is attached to the nitrogen and hydrogen end of HCN, respectively, for ortho- and para-H_{2} complexes and the Σ symmetry has been confirmed for both species. We also reported the MMW spectroscopy of j=1-0 internal rotation band of H_{2}-HCN in 2006 ^{b}, where j is the quantum number for the HCN internal rotation. Although we assigned most of intense lines to the Σ_{1} − Σ_{0} and Π_{1} − Σ_{0} bands of (ortho)H_{2}-HCN, some intense lines are unidentified. To confirm their assignments, we performed the MMW-MMW double resonance spectroscopy in the present study and came to the conclusion that our previous assignments of Σ_{1} − Σ_{0} and Π_{1} − Σ_{0} bands should be changed, and then all of the intense lines are finally assigned to the Π_{1} − Σ_{0} (R_{0}, R_{1},R_{2}, Q_{1}, Q_{2}, and P_{2}) and Σ_{1} − Σ_{0} (R_{0}, and P_{2}) bands. The band origins of the Σ_{1}−Σ_{0} and Π_{1}−Σ_{0} bands of (ortho)H_{2}-HCN newly determined are 187 and 165 GHz, respectively. They are larger than those of Ne-HCN (133 and 107 GHz) but comparable with those of Ar-HCN (165 and 182 GHz, their order is reversed) indicating that the potential anisotropy of (ortho)H_{2}-HCN is larger than that of Ne-HCN but comparable with that of Ar-HCN. The mean square amplitudes of HCN for excited states ( 57^{∘} and 51^{∘} for Σ_{1} and Π_{1}), given by the analysis of hyperfine structure of the nitrogen nucleus, are much larger than that (33^{∘}) of the ground Σ_{0} state. A plenty of weak lines in the 100-300 GHz region are still unassigned, possibly due to the higher internal rotation bands of ortho-H_{2} complex, such as the Δ_{1} − Π_{0} band, as well as the fundamental bands (Σ_{1} − Σ_{0} and Π_{1} − Σ_{0}) of para-H_{2} complex. Analysis of these weak bands and survey in the region with pure para-H_{2} sample are now in progress.

aM. Ishiguro, T. Tanaka, K. Harada, C. J. Whitham and K. Tanaka, J. Chem. Phys. 115, 5155 (2001).

bM. Hagi, K. Harada, and K. Tanaka, The 61st International Symposium on Molecular Spectroscopy, TE01, (2006).